# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v1

@@ Abstract
Index Terms:
Semiparametric models, elliptically symmetric distributions, nuisance parameters, shape matrix, scatter matrix, efficiency, Fisher information matrix, Cramér-Rao bound.
publicationid: pubid: $00$ 2023 IEEE
@@ I Introduction
Asemiparametric model is a statistical model that involves not only a finite-dimensional parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ but also an infinite-dimensional parameter, i.e. a function $g$, that often represents a nuisance parameter. This characterization is general enough to include many well-known examples: the symmetric location model, linear and logistic regression, errors in variables models, missing data and censoring models, copula models and even times series models such as ARMA or ARCH models. We refer the reader to [1, Sect. 4] and [2] for a discussion of how the semiparametric formalism applies to the above-mentioned examples. The clear advantage of adopting a semiparametric model is in the potential gain in term of robustness with rest to some missing knowledge about the random experiment at hand that are indeed required when we use a parametric model. On the other hand, the fact that the function $g$ is left fully unspecified may lead to some efficiency losses in the estimation of the parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ in respect of the parametric case. Nevertheless, there are cases in which parametric and semi-parametric efficiency coincide. In other words, in these specific cases, we can gain robustness without losing efficiency. The simplest example is the symmetric location model, in which the goal is to estimate the location parameter without assuming any specific functional form for the (symmetric) density of the data [1, Sect. 3.4, Ex. 1]. In this work, we will analyze the parametric and the semiparametric efficiency for the estimation of the parameter vector of interest when the considered semiparametric model contains nuisance parameters of both finite and infinite dimension. Specifically, we will focus on the statistical model of the elliptically symmetric distributions.
@@
To make this work as self-contained as possible, in Sect. II we present the geometrical tools (i.e. projections and tagent spaces) that will be at the core of our analysis of the semiparametric elliptical model. Moreover, three fundamental Lemmas are provided and their proofs detailed in the Appendix. We move then to Sect. III where, after a brief recall of the main definitions and properties of the Real Elliptically Symmetric (RES) distributions, we provide an extensive discussion of score vectors and related Fisher Information Matrix (FIM) for the parametric RES model. In doing this, we will make use of some fundamental outcomes obtained in [9, 10] for the inference of the location $\bm{\mu}$ and the shape $\mathbf{V}_{S}$ in the presence of the nuisance scale parameter $s$. Furthermore, we introduce the original analysis of an equivalent RES model, important in many applications, parameterized by $\bm{\mu}$ and the constrained covariance/shape matrix $\bm{\Sigma}$. In Sect. IV the geometrical tools introduced in Sect. II will be specified for the semiparametric RES elliptical models under considerations, while our main results of the first part of the paper is collected in Sect. V and in particular in Proposition V.2, where the closed-form expressions of parametric and semiparametric FIM and the related information bounds for the models in M1, M2 and M3 are provided. Proposition V.2 plays a crucial role in formalizing and fully understanding the relationship between the (lack of) efficiency losses and the finite and infinite-dimensional nuisance terms involved in the considered elliptical parametric and semiparametric models.
@@
So we ask ourselves the same question as before: is it possible to relax the parametric assumption, moving from E1 to E2, without losing efficiency? The answer to this question depends on the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ at hand. Proposition VI.1, which represents the main result of Sect. VI, provides a condition to determine whether the given parameterization leads to efficiency losses or not. As an example, this condition will be applied to two cases, important in many applications, that are i) the parameterization $=(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}\mapsto(\bm{\mu}(\bm{\gamma}),\bm{\Sigma}(\bm{\xi}))$, i.e. when the location and the scatter matrix have no parameters in common and the ii) “low-rank” parameterization. Then, Sect. VII generalizes all the results previously obtained in the context of RES distributions to the case of Circular Complex Elliptically Symmetric (C-CES) and Non Circular CES (NC-CES) distributions. Finally Sect. VIII concludes the paper and the technical proofs are reported in the Appendix.
@@
Notation: Italics indicates scalar quantities ($a$), lower case and upper case boldface indicate column vectors ($\mathbf{a}$) and matrices ($\mathbf{A}$). The superscripts $T$, *, $H$ and ##{\#}# indicate the transpose, the complex conjugate, the Hermitian and the Moore-Penrose inverse operators such that $\mathbf{A}^{-T}\triangleq(\mathbf{A}^{-1})^{T}=(\mathbf{A}^{T})^{-1}$, $\mathbf{A}^{-*}\triangleq(\mathbf{A}^{-1})^{*}=(\mathbf{A}^{*})^{-1}$, $\mathbf{A}^{-H}\triangleq(\mathbf{A}^{-1})^{H}=(\mathbf{A}^{H})^{-1}$ and $\mathbf{A}^{\#T}\triangleq(\mathbf{A}^{\#})^{T}=(\mathbf{A}^{T})^{\#}$. Moreover, $\mathbf{A}^{-1/2}\triangleq(\mathbf{A}^{-1})^{1/2}$ denotes any square root of the inverse of the symmetric positive definite matrix $\mathbf{A}$. Each entry of a vector $\mathbf{a}$ and of a matrix $\mathbf{A}$ is indicated as $a_{i}\triangleq[\mathbf{a}]_{i}$ and $a_{ij}\triangleq[\mathbf{A}]_{ij}$, respectively. The symbol $\mathrm{vec}$ indicates the standard vectorization operator that maps column-wise the entry of an $m\times m$ matrix $\mathbf{A}$ in an $m^{2}$-dimensional column vector $\mathrm{vec}(\mathbf{A})$. The Hadamard product $\mathbf{B}\odot\mathbf{C}$ is the matrix whose $(i,j)$-th element is $[\mathbf{B}]_{ij}[\mathbf{C}]_{ij}$. The Kronecker product ${\mathbf{B}}\otimes{\mathbf{C}}$ denotes the block matrix whose $(i,j)$ block element is $b_{ij}{\mathbf{C}}$, and the commutation matrix ${\mathbf{K}}_{m}$ is such that $\mathrm{vec}(\mathbf{A}^{T})={\bf K}_{m}\mathrm{vec}(\mathbf{A})$. The operator $\underline{\mathrm{vec}}(\mathbf{A})$ defines the $m^{2}-1$-dimensional vector obtained from $\mathrm{vec}(\mathbf{A})$ by deleting its first element, i.e. $\mathrm{vec}(\mathbf{A})\triangleq[a_{11},\underline{\mathrm{vec}}(\mathbf{A})^{T}]^{T}$. For any $m\times m$ symmetric matrix $\mathbf{A}$, $\mathrm{vecs}(\mathbf{A})$ indicates the $m(m+1)/2$-dimensional vector of the entries of the lower triangular part of $\mathbf{A}$. The duplication matrix $\mathbf{D}_{m}$ is implicitly defined as the unique $m^{2}\times m(m+1)/2$ matrix satisfying $\mathbf{D}_{m}\mathrm{vecs}(\mathbf{A})=\mathrm{vec}(\mathbf{A})$ for any symmetric matrix $\mathbf{A}$ [11, 12]. Let us now implicitly define the operator $\underline{\mathrm{vecs}}(\cdot)$ as $\mathrm{vecs}(\mathbf{A})\triangleq[a_{11},\underline{\mathrm{vecs}}(\mathbf{A})^{T}]^{T}$. The identity matrix of dimension $m$ is indicated as $\mathbf{I}_{m}$ and $\mathbf{e}_{k}$ indicates the $k$th vector of the canonical basis of $\mathbb{R}^{m^{2}}$. Moreover, let $\underline{\mathbf{I}}_{m}$ be the operator such that $\underline{\mathrm{vecs}}(\mathbf{A})=\underline{\mathbf{I}}_{m}\mathrm{vecs}(\mathbf{A})$ that can be obtained from $\mathbf{I}_{m(m+1)/2}$ by removing its first row.
@@
[…]
Let us now introduce the Hilbert space $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}})$ as the (infinite-dimensional) linear space of the $\mathfrak{B}(\mathcal{X})$-measurable scalar functions with zero-mean and finite variance:
$\mathcal{H}\triangleq\left\{f:\mathcal{X}\rightarrow\mathbb{R}|{E_{0}\{h\}}=0,{E_{0}\{h^{2}\}}<+\infty\right\},$ (1)
@@
[…]
We note that the norm associated to the inner product in (2) is $||h||_{\mathcal{H}}=\sqrt{{E_{0}\{h^{2}\}}}$ that is the standard deviation of $h\in\mathcal{H}$.
Let us now introduce the $q$-replicating Hilbert space $\mathcal{H}_{q}=\mathcal{H}\times\cdots\times\mathcal{H}$ as the linear space of the $\mathfrak{F}$-measurable, $q$-variate functions $\mathbf{h}:\mathcal{X}\rightarrow\mathbb{R}^{q}$. This set can clearly be obtained as $q$ Cartesian products of $\mathcal{H}$ in (1) and this explain the name of $q$-replicating space.
Remark: It is worth noticing that the $q$-replicating space $\mathcal{H}_{q}$ has been introduced only to simplify some notation and to improve the readability of some result proposed in the following sections. However, it does not introduce any additional geometrical structure since it is just composed of $q$ copies of $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}})$. In particular, when we work on (subspaces of) $\mathcal{H}$, instead of $\mathcal{H}_{q}$, all the operators, as expectations and projections, has to be interpreted component-wise (see e.g. [13, Sect. 2.4, Remark 2] or [1, Sect. 2.4]) as it is proved in Lemma .1 and Theorem .1 reported in Appendix .1. Specifically, for each $\mathbf{h}=(h_{1},\ldots,h_{q})^{T}\in\mathcal{H}_{q}$ s.t. $h_{i}\in\mathcal{H}$ and $u\in\mathcal{U}\subseteq\mathcal{H}$, we have:
$[\Pi(\mathbf{h}|\mathcal{U})]_{i}\triangleq\Pi([\mathbf{h}]_{i}|\mathcal{U})=\Pi(h_{i}|\mathcal{U}),$ (3)
$[{E_{0}\{\mathbf{h}u\}}]_{i}\triangleq{E_{0}\{[\mathbf{h}]_{i}u\}}=\left\langle h_{i},u\right\rangle_{\mathcal{H}},$ (4)
@@
II Models, tangent spaces and score vectors
A semiparametric model is a set of probability density function (pdf) parameterized by a finite dimensional vector of interest and by an infinite-dimensional parameter, i.e. a function, that generally plays the role of a nuisance term. The study of the estimation efficiency in semiparametric models, along with the derivation of the relevant information bounds, is a well-established topic in the statistical literature (see e.g. [1, 13, 14] and the reference therein). We refer the reader to [15, 16] for a tutorial introduction on this subject. In this work, we focus on the case in which the considered semiparametric model involves an additional finite dimensional nuisance vector, along with the infinite-dimensional nuisance term. More formally, let us consider the general semiparametric model:
@@ II Models, tangent spaces and score vectors
[…]
@@
for $j=1,\ldots,r$. Under the regularity conditions discussed in [17, Sects. 6.2, 6.3], it is immediate to verify that $[\mathbf{s}_{\bm{\gamma}_{0}}]_{i},[\mathbf{s}_{\bm{\xi}_{0}}]_{j}\in\mathcal{H}$, $\forall i,j$, i.e. they have zero-mean and finite variance.
We can now introduce the (finite-dimensional) tangent space of the parametric submodels $\mathcal{P}_{2}$ as the linear span of $\mathbf{s}_{\bm{\xi}_{0}}$ in $\mathcal{H}$ ([13, Sect. 2.3], [16, eq. (5.24)]):
@@
According to the previous definition, let us define the efficient score vector $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ for the vector $\bm{\gamma}_{0}$ of the parameter of interest in the parametric submodel $\mathcal{P}_{1,2}$ in (8) as ([13, Sect. 3.4], [1, Sect. 2.4] and [16, Def. 4]):
@@
is the FIM for $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{2}$ in (7). Then, we have that $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ can be explicitly expressed as:
@@
Moreover, it can be noted that $\mathcal{T}_{2}+\mathcal{T}_{3}$ can be expressed as the (direct) sum of the two orthogonal subpsaces $\mathcal{T}_{2}$ and $((\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\bot})$. Then, from (20), we immediately have that:
@@
for the parameter of interest $\bm{\gamma}_{0}$ in the presence of both the finite- and infinite dimensional nuisance parameters $\bm{\xi}_{0}$ and $g_{0}$ with the efficient FIM
@@
for $\bm{\gamma}_{0}$ in the presence of only the finite-dimensional nuisance $\bm{\xi}_{0}$, while $g_{0}$ is known. Specifically, for this comparison, we will make use of the following lemma, given without proof in [1, Sect. 3.4, Prop. 3] and of which we provide a full proof in the Appendix 153 of this work.
@@
It is worth noticing that the matrix $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ is the efficient FIM for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance vector $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{1,2}$ in (8). Consequently the CRB for $\bm{\gamma}_{0}$ in the presence of $\bm{\xi}_{0}$ is given by:
@@
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\triangleq\ \mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})\perp\mathcal{T}_{3}.$ (28)

### v2

@@ Abstract
publicationid: pubid: 0000–0000 © 2023 IEEE
@@ I Introduction
Asemiparametric model is a statistical model that involves not only a finite-dimensional parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ but also an infinite-dimensional parameter, i.e. a function $g$, that often represents a nuisance parameter. This characterization is general enough to include many well-known examples: the symmetric location model, linear and logistic regression, errors in variables models, missing data and censoring models, copula models and even times series models such as ARMA or ARCH models. We refer the reader to [1, Sect. 4] and [2] for a discussion of how the semiparametric formalism applies to the above-mentioned examples. The clear advantage of adopting a semiparametric model is in the potential gain in term of robustness with respect to some missing knowledge about the random experiment at hand that are indeed required when we use a parametric model. On the other hand, the fact that the function $g$ is left fully unspecified may lead to some efficiency losses in the estimation of the parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ in respect of the parametric case. Nevertheless, there are cases in which parametric and semi-parametric efficiency coincide. In other words, in these specific cases, we can gain robustness without losing efficiency. The simplest example is the symmetric location model, in which the goal is to estimate the location parameter without assuming any specific functional form for the (symmetric) density of the data [1, Sect. 3.4, Ex. 1]. In this work, we will analyze the parametric and the semiparametric efficiency for the estimation of the parameter vector of interest when the considered semiparametric model contains nuisance parameters of both finite and infinite dimension. Specifically, we will focus on the statistical model of the elliptically symmetric distributions.
@@
To make this work as self-contained as possible, in Sect. II we present the geometrical tools (i.e. projections and tagent spaces) that will be at the core of our analysis of the semiparametric elliptical model. Moreover, two fundamental lemmas are provided and their proofs detailed in the Appendix. We move then to Sect. III where, after a brief recall of the main definitions and properties of the Real Elliptically Symmetric (RES) distributions, we provide an extensive discussion of score vectors and related Fisher Information Matrix (FIM) for the parametric RES model. In doing this, we will make use of some fundamental outcomes obtained in [9, 10] for the inference of the location $\bm{\mu}$ and the shape $\mathbf{V}_{S}$ in the presence of the nuisance scale parameter $s$. In Sect. IV the geometrical tools introduced in Sect. II will be specified for the semiparametric RES elliptical models under considerations, while our main results of the first part of the paper is collected in Sect. V and in particular in Proposition V.2, where the closed-form expressions of parametric and semiparametric FIM and the related information bounds for the models in M1, M2 and M3 are provided. Proposition V.2 plays a crucial role in formalizing and fully understanding the relationship between the (lack of) efficiency losses and the finite and infinite-dimensional nuisance terms involved in the considered elliptical parametric and semiparametric models.
@@
So we ask ourselves the same question as before: is it possible to relax the parametric assumption, moving from E1 to E2, without losing efficiency? The answer to this question depends on the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ at hand. Proposition VI.1, which represents the main result of Sect. VI, provides a condition to determine whether the given parameterization leads to efficiency losses or not. As an example, this condition will be applied to two cases, important in many applications, that are i) the parameterization $(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}\mapsto(\bm{\mu}(\bm{\gamma}),\bm{\Sigma}(\bm{\xi}))$, i.e. when the location and the scatter matrix have no parameters in common and the ii) “low-rank” parameterization. Then, Sect. VII generalizes all the results previously obtained in the context of RES distributions to the case of Circular Complex Elliptically Symmetric (C-CES) and Non Circular CES (NC-CES) distributions. Finally Sect. VIII concludes the paper and the technical proofs are reported in the Appendix.
@@
Notation: Italics indicates scalar quantities ($a$), lower case and upper case boldface indicate column vectors ($\mathbf{a}$) and matrices ($\mathbf{A}$). The superscripts $T$, *, $H$ and #{\#} indicate the transpose, the complex conjugate, the Hermitian and the Moore-Penrose inverse operators such that $\mathbf{A}^{-T}\triangleq(\mathbf{A}^{-1})^{T}=(\mathbf{A}^{T})^{-1}$, $\mathbf{A}^{-*}\triangleq(\mathbf{A}^{-1})^{*}=(\mathbf{A}^{*})^{-1}$, $\mathbf{A}^{-H}\triangleq(\mathbf{A}^{-1})^{H}=(\mathbf{A}^{H})^{-1}$ and $\mathbf{A}^{\#T}\triangleq(\mathbf{A}^{\#})^{T}=(\mathbf{A}^{T})^{\#}$. Moreover, $\mathbf{A}^{-1/2}\triangleq(\mathbf{A}^{-1})^{1/2}$ denotes any square root of the inverse of the symmetric positive definite matrix $\mathbf{A}$. Each entry of a vector $\mathbf{a}$ and of a matrix $\mathbf{A}$ is indicated as $a_{i}\triangleq[\mathbf{a}]_{i}$ and $a_{ij}\triangleq[\mathbf{A}]_{ij}$, respectively. The symbol $\mathrm{vec}$ indicates the standard vectorization operator that maps column-wise the entry of an $m\times m$ matrix $\mathbf{A}$ in an $m^{2}$-dimensional column vector $\mathrm{vec}(\mathbf{A})$. The Hadamard product $\mathbf{B}\odot\mathbf{C}$ is the matrix whose $(i,j)$-th element is $\mathbf{B}]_{ij}[\mathbf{C}]_{ij}$. The Kronecker product ${\mathbf{B}}\otimes{\mathbf{C}}$ denotes the block matrix whose $(i,j)$ block element is $b_{ij}{\mathbf{C}}$, and the commutation matrix ${\mathbf{K}}_{m}$ is such that $\mathrm{vec}(\mathbf{A}^{T})={\bf K}_{m}\mathrm{vec}(\mathbf{A})$. The operator $\underline{\mathrm{vec}}(\mathbf{A})$ defines the $m^{2}-1$-dimensional vector obtained from $\mathrm{vec}(\mathbf{A})$ by deleting its first element, i.e. $\mathrm{vec}(\mathbf{A})\triangleq[a_{11},\underline{\mathrm{vec}}(\mathbf{A})^{T}]^{T}$. For any $m\times m$ symmetric matrix $\mathbf{A}$, $\mathrm{vecs}(\mathbf{A})$ indicates the $m(m+1)/2$-dimensional vector of the entries of the lower triangular part of $\mathbf{A}$. The duplication matrix $\mathbf{D}_{m}$ is implicitly defined as the unique $m^{2}\times m(m+1)/2$ matrix satisfying $\mathbf{D}_{m}\mathrm{vecs}(\mathbf{A})=\mathrm{vec}(\mathbf{A})$ for any symmetric matrix $\mathbf{A}$ [11, 12]. Let us now implicitly define the operator $\underline{\mathrm{vecs}}(\cdot)$ as $\mathrm{vecs}(\mathbf{A})\triangleq[a_{11},\underline{\mathrm{vecs}}(\mathbf{A})^{T}]^{T}$. The identity matrix of dimension $m$ is indicated as $\mathbf{I}_{m}$ and $\mathbf{e}_{k,l}$ indicates the $k$th vector of the canonical basis of $\mathbb{R}^{l}$. Moreover, let $\underline{\mathbf{I}}_{m}$ be the operator such that $\underline{\mathrm{vecs}}(\mathbf{A})=\underline{\mathbf{I}}_{m}\mathrm{vecs}(\mathbf{A})$ that can be obtained from $\mathbf{I}_{m(m+1)/2}$ by removing its first row.
@@
Let $(\mathcal{X},\mathfrak{B}(\mathcal{X}),P_{0})$ be a probability space where the sample space $\mathcal{X}$ is a subset of $\mathbb{R}^{m}$, $\mathfrak{B}(\mathcal{X})$ is the Borel $\sigma$-algebra on $\mathcal{X}$ and $P_{0}$ is a probability measure. Moreover, $P_{0}$ is assumed to be absolutely continuous with probability density function (pdf), associated to the Lebesgue measure on $\mathbb{R}^{m}$, given by $dP_{0}(\mathbf{x})=p_{0}(\mathbf{x})d\mathbf{x}$. Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be an $\mathfrak{B}(\mathcal{X})$-measurable function, then $E_{0}\{f\}\triangleq\int f(\mathbf{x})dP_{0}(\mathbf{x})$ indicates its expectation w.r.t. $P_{0}$. For random variables or vectors, $_{d}$ stands for ”has the same distribution as”.
Let us now introduce the Hilbert space (,,)($\mathcal{H}$$\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$) as the (infinite-dimensional) linear space of the $\mathfrak{B}(\mathcal{X})$-measurable scalar functions with zero-mean and finite variance:
$\mathcal{H}\triangleq\left\{h:\mathcal{X}\rightarrow\mathbb{R}|{E_{0}\{h\}}=0,{E_{0}\{h^{2}\}}<+\infty\right\},$ (1)
@@
$\left\langle h_{1},h_{2}\right\rangle_{\mathcal{H}}\triangleq{E_{0}\{h_{1}h_{2}\}}=\int_{\mathcal{X}}h_{1}(\mathbf{x})h_{2}(\mathbf{x})dP_{0}(\mathbf{x}),\quad\forall h_{1},h_{2}\in\mathcal{H}.$ (2)
We note that the norm associated to the inner product in (2) is $h||_{\mathcal{H}}=\sqrt{{E_{0}\{h^{2}\}}}$ that is the standard deviation of $h\in\mathcal{H}$.
Let us now introduce the $q$-replicating Hilbert space $\mathcal{H}_{q}=\mathcal{H}\times\cdots\times\mathcal{H}$ as the linear space of the $\mathfrak{B}(\mathcal{X})$-measurable, $q$-variate functions $\mathbf{h}:\mathcal{X}\rightarrow\mathbb{R}^{q}$. This set can clearly be obtained as $q$ Cartesian products of $\mathcal{H}$ in (1) and this explain the name of $q$-replicating space.
Remark: It is worth noticing that the $q$-replicating space $\mathcal{H}_{q}$ has been introduced only to simplify some notation and to improve the readability of some result proposed in the following sections. However, it does not introduce any additional geometrical structure since it is just composed of $q$ copies of (,,)($\mathcal{H}$$\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$). In particular, when we work on (subspaces of) $\mathcal{H}$, instead of $\mathcal{H}_{q}$, all the operators, as expectations and projections, has to be interpreted component-wise (see e.g. [13, Sect. 2.4, Remark 2] or [1, Sect. 2.4]) as it is proved in Lemma .1 and Theorem .1 reported in Appendix .1. Specifically, for each $\mathbf{h}=(h_{1},\ldots,h_{q})^{T}\in\mathcal{H}_{q}$ s.t. $h_{i}\in\mathcal{H}$ and $u\in\mathcal{U}\subseteq\mathcal{H}$, we have:
$\Pi(\mathbf{h}|\mathcal{U})]_{i}\triangleq\Pi([\mathbf{h}]_{i}|\mathcal{U})=\Pi(h_{i}|\mathcal{U}),$ (3)
${E_{0}\{\mathbf{h}u\}}]_{i}\triangleq{E_{0}\{[\mathbf{h}]_{i}u\}}=\left\langle h_{i},u\right\rangle_{\mathcal{H}},$ (4)
@@
II Models, score vectors and tangent spaces
A semiparametric model is a set of probability density functions (pdf) parameterized by a finite dimensional vector of interest and by an infinite-dimensional parameter, i.e. a function, that generally plays the role of a nuisance term. The study of the estimation efficiency in semiparametric models, along with the derivation of the relevant information bounds, is a well-established topic in the statistical literature (see e.g. [1, 13, 14] and the reference therein). We refer the reader to [15, 16] for a tutorial introduction on this subject. In this work, we focus on the case in which the considered semiparametric model involves an additional finite-dimensional nuisance vector, along with the infinite-dimensional nuisance term. More formally, let us consider the general semiparametric model:
@@ II Models, score vectors and tangent spaces
where $\Gamma\subseteq\mathbb{R}^{q}$ is the set of the (finite-dimensional) parameter vectors $\bm{\gamma}$ of interests, $\Psi\subseteq\mathbb{R}^{r}$ is the set of (finite-dimensional) nuisance parameter vectors $\bm{\xi}$ and $\mathcal{G}$ is the set of the (infinite-dimensional) nuisance functions $g$. We indicate as $\bm{\gamma}_{0}\in\Gamma$, $\bm{\xi}_{0}\in\Psi$ and $g_{0}\in\mathcal{G}$ the true, but unknown, related quantities and with $E_{0}\{f\}\triangleq\int f(\mathbf{x})dP_{X}(\mathbf{x}|\bm{\gamma}_{0},\bm{\xi}_{0},g_{0})$ the expectation of a $\mathfrak{B}(\mathcal{X})$-measurable function $f$ w.r.t. the true distribution $P_{0}(\mathbf{x})=P_{X}(\mathbf{x}|\bm{\gamma}_{0},\bm{\xi}_{0},g_{0})$.
@@
for $j=1,\ldots,r$. Under the regularity conditions discussed in [17, Sects. 6.2, 6.3], it is immediate to verify that $\mathbf{s}_{\bm{\gamma}_{0}}]_{i},[\mathbf{s}_{\bm{\xi}_{0}}]_{j}\in\mathcal{H}$, $\forall i,j$, i.e. they have zero-mean and finite variance.
We can now introduce the (finite-dimensional) tangent space of the parametric submodels $\mathcal{P}_{2}$ as the linear span of $\mathbf{s}_{\bm{\xi}_{0}}$ in $\mathcal{H}$ [13, Sect. 2.3], [16, eq. (5.24)]:
@@
According to the previous definition, let us define the efficient score vector $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ for the vector $\bm{\gamma}_{0}$ of the parameter of interest in the parametric submodel $\mathcal{P}_{1,2}$ in (8) as [13, Sect. 3.4], [1, Sect. 2.4] and [16, Def. 4]:
@@
is the Fisher Information Matrix (FIM) for $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{2}$ in (7). Then, we have that $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ can be explicitly expressed as:
@@
Moreover, it can be noted that $\mathcal{T}_{2}+\mathcal{T}_{3}$ can be expressed as the direct sum of the two orthogonal subpsaces $\mathcal{T}_{2}$ and $((\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\bot})$. Then, from (20), we immediately have that:
@@
for the parameter of interest $\bm{\gamma}_{0}$ in the presence of both the finite- and infinite-dimensional nuisance parameters $\bm{\xi}_{0}$ and $g_{0}$ with the efficient FIM
@@
for $\bm{\gamma}_{0}$ in the presence of only the finite-dimensional nuisance $\bm{\xi}_{0}$, while $g_{0}$ is known. Specifically, for this comparison, we will make use of the following lemma, given without proof in [1, Sect. 3.4, Prop. 3] and of which we provide a full proof in the Appendix .2 of this work.
@@
It is worth noticing that the matrix $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ is the efficient FIM for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance vector $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{1,2}$ in (8). Consequently the Cramér-Rao bound (CRB) for $\bm{\gamma}_{0}$ in the presence of $\bm{\xi}_{0}$ is given by:
@@
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})\triangleq\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\perp\mathcal{T}_{3}.$ (28)